OpenStudy (anonymous):

MEDAL AND FAN The function f(x) is shown below: f(x)=(4)x+5 The function f(x) is shifted to the right by 13 units. Which of the following best represents the new function? A) f(x)=(4)x-8 B) f(x)=(4)x+13 (my choice) C) f(x)=(4)x+18 D) f(x)=4x-13

OpenStudy (anonymous):

@Nnesha Could you please check this?

OpenStudy (anonymous):

anyone?

Nnesha (nnesha):

they said right so you know its opposite - axis ---->right side +axis -------> left side

OpenStudy (anonymous):

Oh so it would be D?

OpenStudy (anonymous):

Because than it would be opposite

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

@Nnesha

OpenStudy (solomonzelman):

\(\large\color{black}{f(x)=4x-8}\) I'll try to give you a couple of examples of shifts, hope they can help.... \(\large\color{ blue }{\large {\bbox[5pt, lightyellow ,border:2px solid white ]{ \large\text{ }\\ \begin{array}{|c|c|c|c|} \hline \texttt{Shifts} ~~~\tt from~~~ {f(x)~~~\tt to~~~g(x)}&~\tt{c~~~units~~~~} \\ \hline \\f(x)= \sqrt[4]{x} ~~~~\rm{\Rightarrow}~~~~ g(x)= \sqrt[4]{x \normalsize\color{red }{ -~\rm{c}} } &~\rm{to~~the~~right~} \\ \text{ } \\ f(x)= \sqrt[4]{x} ~~~~\rm{\Rightarrow}~~~~ g(x)= \sqrt[4]{x \normalsize\color{red}{ +~\rm{c}} } &~\rm{to~~the~~left ~} \\ \text{ } \\ f(x)= \sqrt[4]{x} ~~~~\rm{\Rightarrow}~~~~ g(x)= \sqrt[4]{x} \normalsize\color{red}{ +~\rm{c} } &~\rm{up~} \\ \text{ } \\ f(x)= \sqrt[4]{x} ~~~~\rm{\Rightarrow}~~~~ g(x)= \sqrt[4]{x} \normalsize\color{red}{ -~\rm{c} } &~\rm{down~} \\ \\ \hline \end{array} }}}\) \(\large\color{ teal }{\large {\bbox[5pt, lightcyan ,border:2px solid white ]{ \large\text{ }\\ \begin{array}{|c|c|c|c|} \hline \texttt{Shifts} ~~~\tt from~~~ {f(x)~~~\tt to~~~g(x)}&~\tt{c~~~units~~~~} \\ \hline \\f(x)= x^2 ~~~~~\rm{\Rightarrow}~~~~ g(x)= (x \normalsize\color{red}{ -~\rm{c} })^2 &~\rm{to~~the~~right~} \\ \text{ } \\ f(x)= x^2 ~~~~~\rm{\Rightarrow}~~~~ g(x)= (x \normalsize\color{red}{ +~\rm{c} })^2&~\rm{to~~the~~left ~} \\ \text{ } \\ f(x)= x^2 ~~~~~\rm{\Rightarrow}~~~~ g(x)= x^2 \normalsize\color{red}{ +~\rm{c} } &~\rm{up~} \\ \text{ } \\ f(x)= x^2 ~~~~~\rm{\Rightarrow}~~~~ g(x)= x^2 \normalsize\color{red}{ -~\rm{c} } &~\rm{down~} \\ \\ \hline \end{array} }}}\)

Nnesha (nnesha):

yes right sorry i lost te net connection so.

OpenStudy (solomonzelman):

me too, lol

Nnesha (nnesha):

:P

OpenStudy (anonymous):

:D thank you both! </3 I can't give two medals though

OpenStudy (solomonzelman):

ha ha

OpenStudy (solomonzelman):

I don't need any medals.

OpenStudy (solomonzelman):

take the knowledge....

OpenStudy (anonymous):

Okay :D i still fan u

OpenStudy (solomonzelman):

If you want to, I would certainly not mind. Not that I wouldn't help you only knowing that you would fan me... so ty